# DIVER: A Multi-Stage Approach for Reasoning-intensive Information Retrieval

### v4

@@ 3.2.1 Comparison with Different Retrievers
We compare our DIVER-Retriever with a range of retrieval baselines: (1) the sparse retriever BM25 (2009_bm25); (2) open-source dense retrievers including SBERT (all-mpnet-base-v2, 19emnlp_sBERT), gte-Qwen1.5-7B (gte_qwen1.5), and Qwen3-4B (Qwen3-Embedding-4B, qwen3emb); (3) proprietary models from OpenAI [^9] https://openai.com/index/new-embedding-models-and-api-updates/. (text-embedding-3-large) and Google (text-embedding-preview0409, 24google_gecko); and (4) reasoning-aware retrievers including ReasonIR-8B (25_ReasonIR), RaDeR-7B (25_RaDeR), and Seed1.5-Embedding [^10] https://seed1-5-embedding.github.io/.
@@ 3.2.2 Comparison with Different Expanded Queries
This section compares DIVER-QExpand with several representative zero-shot query expansion methods: (1) GPT-4 Reason-query (25ICLR_BRIGHT), which uses chain-of-thought reasoning steps from GPT-4 as queries, provided in BRIGHT; (2) XRR2111111https://github.com/jataware/XRR2/tree/main, which applies carefully crafted prompts with GPT-4o for query expansion; (3) TongSearch-QR-7B (25_tongsearch_qr), a small-size language model trained by reinforcement learning (RL) for reasoning-based query rewriting; and (4) ThinkQE-14B (25_ThinkQE), a thinking-based query expansion method that iteratively refines queries using feedback from retrieved documents.
@@ 3.2.3 Effect of DIVER-DChunk
As shown in Table 5, the improvement for DIVER-Retriever is consistent across most domains, with larger gains in Psychology (+2.3) and Stack Overflow (+1.2). This suggests that cleaner and semantically aligned chunks provide richer context for dense retrieval models. The impact on BM25 is minimal, which is expected since it is less sensitive to chunk coherence. Our observation aligns with recent work (25_bright+) that also highlights the limitations of BRIGHT’s original chunking and proposes BRIGHT+, a multi-agent LLM-based cleaning and rechunking pipeline to build a higher-quality corpus.

### v5

@@ 3.2.1 Comparison with Different Retrievers
We compare our DIVER-Retriever with a range of retrieval baselines: (1) the sparse retriever BM25 (Robertson and Zaragoza, 2009); (2) open-source dense retrievers including SBERT (all-mpnet-base-v2, Reimers and Gurevych (2019)), gte-Qwen1.5-7B (Li et al., 2023), and Qwen3-4B (Qwen3-Embedding-4B, Zhang et al. (2025b)); (3) proprietary models from OpenAI [^9] https://openai.com/index/new-embedding-models-and-api-updates/. (text-embedding-3-large) and Google (text-embedding-preview0409, Lee et al. (2024)); and (4) reasoning-aware retrievers including ReasonIR-8B (Shao et al., 2025), RaDeR-7B (Das et al., 2025), and Seed1.5-Embedding [^10] https://seed1-5-embedding.github.io/.
@@ 3.2.2 Comparison with Different Expanded Queries
This section compares DIVER-QExpand with several representative zero-shot query expansion methods: (1) GPT-4 Reason-query (Su et al., 2025), which uses chain-of-thought reasoning steps from GPT-4 as queries, provided in BRIGHT; (2) XRR2111111https://github.com/jataware/XRR2/tree/main, which applies carefully crafted prompts with GPT-4o for query expansion; (3) TongSearch-QR-7B (Qin et al., 2025), a small-size language model trained by reinforcement learning (RL) for reasoning-based query rewriting; and (4) ThinkQE-14B (Lei et al., 2025), a thinking-based query expansion method that iteratively refines queries using feedback from retrieved documents.
@@ 3.2.3 Effect of DIVER-DChunk
As shown in Table 5, the improvement for DIVER-Retriever is consistent across most domains, with larger gains in Psychology (+2.3) and Stack Overflow (+1.2). This suggests that cleaner and semantically aligned chunks provide richer context for dense retrieval models. The impact on BM25 is minimal, which is expected since it is less sensitive to chunk coherence. Our observation aligns with recent work (Chen et al., 2025b) that also highlights the limitations of BRIGHT’s original chunking and proposes BRIGHT+, a multi-agent LLM-based cleaning and rechunking pipeline to build a higher-quality corpus.